OpenStudy (anonymous):

I need some math help plz Solve the following system of equations: x+5y=7 2x+4y=-4

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

Solve x+5y=7 for x: x+5y+−5y=7+−5y(Add -5y to both sides) x=−5y+7 Now we have x=−5y+7 <---equation 1 2x+4y=-4 <---equation 2 Plug equation 1 into equation 2 since we know the value of x 2(−5y+7)+4y=−4 solve for y

OpenStudy (anonymous):

wait soo this is the answer?

pooja195 (pooja195):

No you need to find it

OpenStudy (anonymous):

ohhhh i seee

OpenStudy (anonymous):

@pooja195 do you know the steps for solving a system of equations using the substitution method